# Massive scalar perturbations in Kerr Black Holes: near extremal analysis

### v2

@@ VII Exceptional point and hysteresis
Continuing the survey, if we start at the origin of the parameter space, first increase $a/M$, while keeping $M\mu=0$ fixed, and then increase $M\mu$ while keeping $a/M$ fixed, we observe the emergence of unstable modes. In fact, for any spin parameter, if the scalar field is sufficiently heavy, the fundamental mode is no longer stable since the imaginary part of its frequency becomes positive [44]. We have identified a curve, shown in Fig. 10, along which the fundamental QNM frequencies become purely real. Hence, parameters corresponding to points above the curve allow the existence of unstable modes that satisfy the QNM boundary conditions.
Figure 9: The values $a/M$ as a function of $M\mu$ where the imaginary parts of the longest-living mode and the first overtone are equal, defining thus the “coexistence line”. The line begins at the bifurcation point $\{(a/M)_{c},(M\mu)_{c}\}$ and joins the instability line at the extremal limit $a/M=1$.
Figure 10: Scalar field masses $M\mu$ as a function of spin $a/M$ for which the fundamental $m=\ell=1$ QNM is infinitely long-lived. The corresponding frequencies are purely real. Above the curve, unstable modes that satisfy the QNM boundary condition exist.
Figure 11: The representation of the longest living mode as a function of the parameters. The gray (green) area represents the region of linear stability, and the dashed line represent the line of coexistence where the imaginary parts of the longest living mode and the first overtone are equal. The line ends at the bifurcation point $\{(a/M)_{c},(M\mu)_{c}\}$.
Figure 12: A scheme of the path taken to test for hysteresis. The path $ABCD$ encircles $\{(a/M)_{c},(M\mu)_{c}\}$. We test the existence of an exceptional point by testing whether the adiabatic change of parameters from $A$ to $D$ depends on the path taken.
@@
In Fig. 11 we display both curves which were shown separately in Figs. 9 and 10. The green region corresponds to parameters for which all $\ell=m=1$ modes satisfying the QNM boundary conditions are stable. The dashed line is the curve along which $\mathrm{Im}(M\omega_{0})=\mathrm{Im}(M\omega_{1})$. It meets the curve along which the fundamental mode is infinitely-long lived at $M\mu\simeq 0.870750$ and $a/M\rightarrow 1$. More importantly, the other endpoint of the dashed curve in Fig. 11, represented by a yellow star, is exactly the critical point identified in Sec. IV.2. For visualization purposes, the dashed curve and the critical spin parameter $(a/M)_{c}$ have been intentionally plotted out of scale.
@@
The existence of this non-trivial holonomy in parameter space also gives support to the idea that the Riemann surface in which the eigenvalues of the Hamiltonian exist is in fact a two-sheeted cover of $\{(a/M),(M\mu)\}$. We illustrate the behavior in Fig. 13, where we plot $\mathrm{Im}(M\omega)$ for the two dominant QNMs when $\ell=m=1$. We also plot two paths in the three-dimensional space $\{(a/M),(M\mu),-\mathrm{Im}(M\omega)\}$: the solid white curve encircles the exceptional point (like the path $ABCD$ in Fig. 12), while the dashed red curve crosses the coexistence line (like the path $AD$ in Fig. 12), going continuously from one sheet into the other. In other words, each sheet of the covering space corresponds to a different path taken. If we circle around the exceptional point twice, we return both eigenvalues to their initial configuration, thus establishing a two-sheet cover.
@@ VIII Discussion
We anticipate that other black hole systems in which bifurcation points arise [39, 40] will also display exceptional points as two-sheeted branching points of their parameter spaces. Exceptional points and their associated non-trivial holonomy have thus far been linked to phase transitions and to quantum mechanics of open systems [101], impacting the fields of Condensed Matter and Optics. Similarly, we expect that our result regarding the presence of these features in gravitational physics will significantly influence the study of linear perturbations in black holes.
Note added. After completing this work, we became aware of a recent preprint [102], which interprets an avoided crossing near a degeneracy point as the cause of resonances between gravitational $\ell=m=2$ QNMs.

### v3

@@ VII Exceptional point and hysteresis
Continuing the survey, if we start at the origin of the parameter space, first increase $a/M$, while keeping $M\mu=0$ fixed, and then increase $M\mu$ while keeping $a/M$ fixed, we observe a curve where the real part of the parameter $\alpha$, defined in (7), vanishes. As shown in Fig. 10, we find that moving adiabatically the fundamental mode to increasing values of $M\mu$ leads to the line $\mathrm{Re}\,\alpha=0$, which we refer to as the limiting line. Along this line, the imaginary part of the QNM frequency also vanishes, bringing up the question of existence and stability of QNMs beyond it [44]. In fact, beyond the limiting line the real part of $\alpha$ should be negative for the fundamental mode, meaning that the boundary conditions for QNMs, defined in (8), are no longer expressed in terms of the monodromy parameters by Eq. (17).
Figure 9: The values $a/M$ as a function of $M\mu$ where the imaginary parts of the longest-living mode and the first overtone are equal, defining thus the “coexistence line”. The line begins at the bifurcation point $\{(a/M)_{c},(M\mu)_{c}\}$ and joins the limiting line of our analysis – where $\mathrm{Re}\,\alpha=0$ – at the extremal limit $a/M=1$.
Figure 10: The limiting line where $\mathrm{Re}\,\alpha=0$, above which the relation between the QNM boundary conditions and the monodromy parameters given by (17) is no longer valid. On this line the imaginary part of the QNM fundamental mode vanishes, raising the question of existence and instability of QNMs above it.
Figure 11: The representation of the longest living mode as a function of the parameters. The gray (green) region indicates where $\mathrm{Re}\,\alpha>0$, making our analysis of the monodromy parameters valid within this area. The dashed line represents the line of coexistence where the imaginary parts of the longest living mode and the first overtone are equal. The line ends at the bifurcation point $\{(a/M)_{c},(M\mu)_{c}\}$.
Figure 12: A scheme of the path taken to test for hysteresis. The path $ABCD$ encircles $\{(a/M)_{c},(M\mu)_{c}\}$. We verify the existence of an exceptional point by testing whether the adiabatic change of parameters from $A$ to $D$ depends on the path taken.
@@
In Fig. 11 we display both curves which were shown separately in Figs. 9 and 10. The green region represents the portion of the parameter space analyzed in this study with respect to the $\ell=m=1$ QNMs. The dashed line is the curve along which $\mathrm{Im}(M\omega_{0})=\mathrm{Im}(M\omega_{1})$. It meets the limiting curve along which the fundamental mode is infinitely-long lived at $M\mu\simeq 0.870750$ and $a/M\rightarrow 1$. More importantly, the other endpoint of the dashed curve in Fig. 11, represented by a yellow star, is exactly the critical point identified in Sec. IV.2. For visualization purposes, the dashed curve and the critical spin parameter $(a/M)_{c}$ have been intentionally plotted out of scale.
@@
The existence of this non-trivial holonomy in parameter space also gives support to the idea that the surface in which the eigenvalues of the Hamiltonian exist is in fact a two-sheeted cover of $\{(a/M),(M\mu)\}$. We illustrate the behavior in Fig. 13, where we plot $\mathrm{Im}(M\omega)$ for the two dominant QNMs when $\ell=m=1$. We also plot two paths in the three-dimensional space $\{(a/M),(M\mu),-\mathrm{Im}(M\omega)\}$: the solid white curve encircles the exceptional point (like the path $ABCD$ in Fig. 12), while the dashed red curve crosses the coexistence line (like the path $AD$ in Fig. 12), going continuously from one sheet into the other. In other words, each sheet of the covering space corresponds to a different path taken. If we circle around the exceptional point twice, we return both eigenvalues to their initial configuration, thus establishing a two-sheet cover.
In this paper, we have focused our analysis on the lowest-lying QNMs, as ordered by the imaginary part of their frequencies in the massless $M\mu=0$, Schwarzschild $a/M=0$ case. However, the ingredients for exceptional points and hysteresis, as discussed in the preceding paragraphs, may also be present for other pairs of overtones, raising the natural question of whether additional exceptional points exist.
While a thorough analysis lies outside the scope of this article, a preliminary survey found another exceptional point at
$a/M\simeq 0.999854,\qquad M\mu\simeq 0.3191,$ (68)
where the $n=1$ and $n=2$ QNMs for $\ell=m=1$ are degenerate. We have also found that a circuit around this point interchanges the $n=1$ and $n=2$ levels in a manner similar to to the results shown in (66) and (67). Although the analysis is complicated by the near-extremal regime, we may find more exceptional points, involving higher excited states, when $a/M\rightarrow 1$. We will return to the identification and characterization of additional exceptional points in future work.
@@ VIII Discussion
We remark that our article is the first to discuss the role of the mass of a perturbing field in the context of ZDMs and DMs. Building on the eikonal limit arguments presented in Refs. [46, 47], a possible explanation for the mass-induced transition observed between ZDMs and DMs lies in the influence of the mass parameter on the peak of the potential barrier surrounding the black hole. Furthermore, our research suggests a potential connection between exceptional points and transitions from ZDMs to DMs. We believe that our work will contribute to a deeper understanding of ZDM-DM transitions, particularly those associated with QNM degeneracies.
Finally, we anticipate that other black hole systems in which bifurcation points arise [39, 40] will also display exceptional points as two-sheeted branching points of their parameter spaces. Exceptional points and their associated non-trivial holonomy have thus far been linked to phase transitions and to quantum mechanics of open systems [102], impacting the fields of Condensed Matter and Optics. Similarly, we expect that our result regarding the presence of these features in gravitational physics will significantly influence the study of linear perturbations in black holes.
Note added. After completing this work, we became aware of a recent preprint [103], which interprets an avoided crossing [104, 105, 106] near a degeneracy point as the cause of resonances between gravitational $\ell=m=2$ QNMs.